OpenStudy (anonymous):

please help asap What is the force of a a bowl with a mass of 3 kg accelerating at a speed of 18 kg/s/s?

jimthompson5910 (jim_thompson5910):

is that a typo? where it says kg/s/s is it supposed to say m/s/s or m/s^2?

OpenStudy (anonymous):

no its not please help

OpenStudy (anonymous):

nevermind

jimthompson5910 (jim_thompson5910):

well honestly the units for the acceleration make no sense. I'm going to assume they meant to write "m/s/s" which is the same as "m/s^2" force = mass * acceleration f = ma f = (3 kg)*(18 m/s^2) f = (3*18) (kg*m/s^2) f = 54 N so the force is 54 newtons. A newton is equivalent to the unit "kg*m/s^2" I'm using newton's second law http://www.physicsclassroom.com/class/newtlaws/Lesson-3/Newton-s-Second-Law